MD=SD=R/2 (Sign are natural)


1
Expert's answer
2021-02-22T11:58:45-0500

Without going to the "\dfrac{R\text{(Range)}}{2}" part,


 Why do you think "MD = SD" ?

That is NOT generally true (although data can be constructed such that MD = SD).


For example: "x = {[-10, 0, 10]}, SD(x) = \sqrt{\dfrac{(10^2 + 0 + 10^2)}{3}} = ~2.72"


(or another answer if you divide by "N-1" instead of "N" )


 "MD(x) = \dfrac{(10 + 0 + 10)}{3} = 6.66..."


"SD(x) \neq MD(x)"


Hence, "MD\neq SD\neq \dfrac{R}{2}"